a car is accelerating forward with an acceleration 9.8 . a pendulum is suspended from its celling .calculate the angle made by the string with vertical in steady state.

Grade:12

1 Answers

jitender
6 years ago
Due to pseudo force Bob will move in oppositeDirection of motion of car.Pseudo force, f = ma = 9.8mNow in equilibrium or steady state condition is achieved at angle ¤.And NET force on Bob is zero.If T is tension in string then ....Tcos¤ = Mg ........1st eqn Tsin¤ = Ma ........2nd eqn Divide 2/1We get ... Tsin¤/Tcos¤ = Ma/MgTan¤ = a/g = 9.8/9.8 = 1Tan¤ = Tan45.And ¤ = 45°So Bob is stationary at 45° with vertical